Find the distance between the two charges, q1 with 36 nC and q2 with 22 nC. The electric potential energy is 9.912 x 11-3 N x m


1
Expert's answer
2022-05-20T04:53:06-0400
"E_p=k\frac{q_1q_2}{r}"

"r=k\frac{q_1q_2}{E_p}\\
=9*10^9*\frac{36*10^{-9}*22*10^{-9}}{9.912*10^{-3}}=7.2*10^{-4}\:\rm m"